OpenStudy (anonymous):

What is the simplified result of the expression -2i(3-4i+8i squared)

terenzreignz (terenzreignz):

Let's kill the ambiguity... is this what it looks like... \[\Large -2i(3-4i+8i^2)\] ?

OpenStudy (anonymous):

Yes I didn't know how to type the square in the formula

terenzreignz (terenzreignz):

Okay, so, first things first, you know that... \[\large i^2 = -1\]right?

OpenStudy (anonymous):

nope

terenzreignz (terenzreignz):

Seriously? :/ Well anyway, i is defined as such.... \[\Large i = \sqrt{-1}\] So it only makes sense that... \[\Large i^2 = -1\]

OpenStudy (anonymous):

I haven't done this kind of math in 30 years...I'm trying to relearn for a certification test. And during my long career I truly did not use this. :)

terenzreignz (terenzreignz):

Understood, sir/madame... so, granted, and accepted that \[\Large i^2 = -1\] We have an occurrence of \(\large i^2\) right here... \[\Large -2i(3-4i+8\color{red}{i^2})\] So that may be replaced by its equivalent (equal) -1... like so... \[\Large -2i(3-4i+8\color{blue}{(-1)})\]

OpenStudy (anonymous):

oh ok

terenzreignz (terenzreignz):

So simplifying yields... \[\Large -2i(3-4i\color{green}{-8})\] These two terms may be combined into one... \[\Large -2i(\color{orange}3-4i\color{orange}{-8})\] \[\Large -2i(\color{blue}{-5}-4i)\] At which point, all that's left is distributing the -2i

terenzreignz (terenzreignz):

Good luck with your certification test :)

OpenStudy (anonymous):

oh that makes so much sense now...thank you!!! Lynn

terenzreignz (terenzreignz):

No problem, Lynn :) The name's Terence, at your service :D

OpenStudy (anonymous):

Thanks again.